Operations Report – May 16th

Crew 299 Operations Report 5-16-2024

SOL: 4

Name of person filing report: Rishabh Pandey (Engineer)

Non-Nominal Systems: Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: Issues with both suits were noted before our mission, both will not be used during our mission.

Rovers Used: None
General notes on rovers:

Summary of Hab operations:
      WATER USE: 70.12 Gallons (septic tank cleaning might have caused this value to double from yesterday)
      Water (static tank): 309 (using effective value from spreadsheet)
      Static tank pipe heater: off
      Static tank heater: off
      Toilet tank emptied: yes

Summary of internet usage: Catch up on Earthly news, email responses, code development, entertainment

Summary of suits and radios: None used

Summary of GreenHab operations: NA, GreenHab out of operation

Summary of ScienceDome operations: Trimmed radish leaves in an attempt to stop a potential bacterial infection. Soil was checked again. Incident with vial breaking, all glass shards were safely secured and disposed.

Summary of RAM operations: Multimeter and Measuring tape used
Summary of observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site